OpenStudy (anonymous):

please helppppp! I'm horrible at math :/ In the below system, solve for y in the first equation. x + 3y = 6 2x – y = 10 one thirdx + 2 negative one thirdx + 6 –x + 2 negative one thirdx + 2

OpenStudy (anonymous):

y = -1/3x + 2 ?

OpenStudy (anonymous):

looks right to me!

OpenStudy (anonymous):

Is your last choice.

OpenStudy (anonymous):

oh! i missed a few steps...or did something wrong i guess. thank you <3

OpenStudy (anonymous):

Use the substitution method to solve the following system of equations. 4x – y = 3 7x – 9y = –2 (1, 1) (6, –3) (6, 1) (1, –3)

OpenStudy (anonymous):

would it be (1,1)?

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

(1,1) is right for that one